OpenStudy (anonymous):

help, is this true? sin(theta)=x-2 the same as (theta)=arcsin(x-2)

OpenStudy (anonymous):

or how do i solve for theta?

myininaya (myininaya):

no not the same sin(theta)=x-2 gives me any solutions to theta theta=arcsin(x-2) gives only one value for theta

OpenStudy (anonymous):

how would i solve it? im doing trig substitution but i got a theta and i dont know what to substitute, i know i have sintheta=x-2

myininaya (myininaya):

sin(theta)=sin(theta+2 pi *n) for integer n right?

OpenStudy (anonymous):

trig sub you only use one value

myininaya (myininaya):

\[\sin(\theta+2 n \pi) =x-2 \]

myininaya (myininaya):

So now solve for theta like you were

OpenStudy (anonymous):

i bet if you post the original problem myininaya would be all over it.

OpenStudy (anonymous):

maybe post the original problem and say what the substitution was then we can go backwards

OpenStudy (anonymous):

im so lost :( when i integrate i get -cos(theta)+(1/4)cos^4(theta)+(1/3)cos^3(theta)+2theta

OpenStudy (anonymous):

\[\int\limits_{1}^{3}x \sqrt{1-(x-2)^2}dx\]

myininaya (myininaya):

i will brb i have food @satellite73

OpenStudy (anonymous):

lord what a pain but that is ok, you have the answer it looks like the last term will just be \[\sin^{-1}(x-2)\] there is nothing you can do about it

OpenStudy (anonymous):

so that is what theta equals?

OpenStudy (anonymous):

fortunately for you this is easy to evaluate at 3 and at 1

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so i should get pi?

OpenStudy (anonymous):

as my final?

OpenStudy (anonymous):

3 gets you \[\sin^{-1}(1)=\frac{\pi}{2}\] and 1 gets you \[\sin^{-1}(-1)=-\frac{\pi}{2}\] so it is not a big deal

OpenStudy (anonymous):

yes subtract and get \[\pi\]

OpenStudy (anonymous):

not sure about the first part though

OpenStudy (anonymous):

actually yes, the whole thing is pi that is it!

OpenStudy (anonymous):

sweet

OpenStudy (anonymous):

i think (although i am not sure) you can get the answer most easily by changing the limits of integration twice

myininaya (myininaya):

ok so i'm back

OpenStudy (anonymous):

first by \[u=x-2\] change to \[\int_{-1}^1\] and then when you make \[\sin(\theta)=u\] change to \[\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\]

OpenStudy (anonymous):

that way you do not have to go backwards once you get your answer. myininaya knows more about this than i do. maybe i am wrong. your answer is right however

OpenStudy (anonymous):

thanks a lot

myininaya (myininaya):

\[\int\limits\limits_{1}^{3}x \sqrt{1-(x-2)^2}dx \] \[\text{ Let } x-2 = \cos(\theta)\] => \[dx=-\sin(\theta) d \theta\] So we have x=1 to x=3 \[\text{We this in terms of } \theta \text{ since we are putting the expression in terms of } \theta\] Ok so we have: \[1-2=\cos(\theta) \text{ to } 3-2=\cos(\theta)\] \[=>-1=\cos(\theta) \text{ to } 1= \cos(\theta)\] \[=>\theta=\pi \text{ to } \theta=0\] So we have \[\int\limits_{\pi}^{0} \sqrt{1-\cos^2(\theta)} (\cos(\theta)+2) (-\sin(\theta)) d \theta \] and by the way you could have gone with the sin sub if you wanted to